OpenStudy (anonymous):

Match the function with its value. f(x) = x2 - 3x + 3 g(x) = 2x2 -1 f(z2)

OpenStudy (anonymous):

these are the answers w3 +1 7 z4 -3z2 +3 17 z2 -3z +3 2w6 -1

OpenStudy (anonymous):

f(z^2)= (z^2)^2-3z^2+3=z^4-3z^2+3

OpenStudy (anonymous):

wat in the world is tht!!!??? lol

OpenStudy (anonymous):

maybe i did not understand your question

OpenStudy (anonymous):

is it \(f(2)\) ???

OpenStudy (anonymous):

its probably not you im dumb at things u r probably rite and i juss dnt understand

OpenStudy (anonymous):

no its f(zsquared)

OpenStudy (anonymous):

Put \(x = z^2\) in f(x): \[f(z^2) = (z^2)^2 - 3(z^2) + 3\] Now solve it..

OpenStudy (anonymous):

Getting @janessia ??

OpenStudy (cwrw238):

if f(x) = x^2 - 3x - 3 then f(z^2) is what ananymas said - you replace the x in the function by z^2 and simplify

OpenStudy (anonymous):

Just replace x by \(z^2\) in f(x)..

OpenStudy (anonymous):

hmmmmmm ohk ohk hold on u guys im still processing all of this